OpenStudy (anonymous):

h9 p1 5 anyone? plss help...........

OpenStudy (anonymous):

Voltage across capacitor = -Q/C=-A/C

OpenStudy (anonymous):

-125.66

OpenStudy (anonymous):

Numerical values may differ from person to person... So i always go with formula...

OpenStudy (anonymous):

I used this formula for voltage: -L*2*pi*f*sin(2*pi*f*t) and got the correct answer....hope it helps

OpenStudy (anonymous):

h9p1 --- e) -15.7079632679 voltage across capacitor